OpenStudy (eduardo):

A student dissolves 12.g of sucrose C12H22O11 in 300.mL of a solvent with a density of 1.01/gmL . The student notices that the volume of the solvent does not change when the sucrose dissolves in it. Calculate the molarity and molality of the student's solution. Round both of your answers to 2 significant digits. molarity = molality =

OpenStudy (eduardo):

molarity(M) = mole of solute/volume of solvent in L molecular weight of sucrose = 342.30 g/mol, moles of sucrose in the given question = 12/342.30 = 3.5*10-2 mol molarity = 3.5*10-2/300*10-3 = 0.12M answer molality(m) = mole of solute/solvent in Kg density(d) = m/v so mass of solvent in Kg (m) = d*v = 1.01*300*1000 = 3.03*105 so molality = 3.5*10-2/3.03*105 = 1.2*10-7mol/Kg answer will this be correct because i don't know if I'm right on molality or everything overall

OpenStudy (sunnnystrong):

I realize this was posted 3 months ago but... Molarity = moles of solute/liter of solution so: C12H22O11 (Table Sugar) has 342.29 g/mol 12g. x (1 mole of sugar/342.29g of sugar) = 0.035 moles of sugar. 0.035 moles of sugar/ .3 L solution = 0.12 Molarity ----- Molality is moles of solute/kg of solvent So: the density of the solution is 1.01g/mL and there are 300 mL 300mL solution (1.01g/mL) = 303 g of solution soluition-solute=solvent mass 303g solution - 12g. of solute (sugar) = 291g of solvent (.291kg of solvent) so: 0.035 Moles of Sugar / .291 kg solvent = 0.12 Molality